The weighs in kilograms of 3 groups of 5 each are shown in below. Is there unusual variation among the groups?

1st member: 1st group is 50, 2nd group is 60, 3rd group is 53

  2nd member: 1st group is 48, 2nd group is 40, 3rd group is 55

 3rd member: 1st group is 55, 2nd group is 50, 3rd group is 40

 4th member: 1st group is 50, 2nd group is 60, 3rd group is 40

5th member: 1st group is 46, 2nd group is 52, 3rd group is 47

 

 


1
Expert's answer
2021-11-16T14:22:28-0500

One-Way ANOVA test

using online calculator https://goodcalculators.com/, we get:


H0: µ1 = µ2 = µ3    ("all 3 population means are equal")

H1: At least one µi different ("at least one of the 3 population means is not equal to the others")

test statistic:

computed as

"F=\frac{MSR}{MSE}"

where MSR = SSR/dfr = the regression mean square,

MSE = SSE/dfe = the mean square error,

SSR = the regression sum of squares,

SSE = the error sum of squares,

SST = the total sum of squares (SST = SSR + SSE),

dfr = the model degrees of freedom (equal to dfr = k - 1),

dfe = the error degrees of freedom (equal to dfe = n - k - 1)

k = the total number of groups (levels of the independent variable)

n = the total number of valid observations,

dfT = the total degrees of freedom (equal to dfT = dfr + dfe = n - 1)


for "\alpha=0.05" :

p-value > 0.05

so, we can conclude that there is no unusual variation among the groups